# Slot State Space Models

### v5

@@ 8 Conclusion & Limitations
SlotSSMs’ success illustrates the importance of designing architectures that align with the problem domain’s underlying modular structure. It also paves the way for future research in modular and object-centric sequence modeling. However, it has some limitations that future studies could address. First, although the proposed architecture is not only applicable to video modeling but also to other modalities like text, we have not explored this aspect in this study. It remains a matter for future work. Second, due to our academic research lab’s computing resource constraints, we were unable to significantly scale up the proposed model to industry-scale in terms of model size and data size. Lastly, future studies should investigate the effect of increased visual complexity in videos.
Impact Statement
The introduction of SlotSSMs, a novel framework that incorporates independent mechanisms into State Space Models (SSMs), has the potential to significantly impact the field of sequence modeling. By leveraging the modular structure inherent in many real-world processes, SlotSSMs offers a more intuitive and effective approach to modeling long-range temporal dependencies in object-centric video understanding and prediction tasks. The substantial performance gains demonstrated by SlotSSMs over existing sequence modeling methods highlight the importance of designing architectures that align with the underlying structure of the problem domain. This breakthrough could lead to the development of more efficient and accurate models for a wide range of applications, such as robotics, autonomous vehicles, and video surveillance systems. Moreover, the success of SlotSSMs in capturing the modular nature of real-world processes could inspire further research into modular and object-centric sequence modeling. This could result in the development of even more advanced architectures that can better handle the complexity and diversity of real-world data. Because this is a general backbone architecture for sequence modeling, it doesn’t raise direct ethical concerns. However, its ethical implications depend on the way downstream application developers use the model.
@@ Acknowledgements
This work is supported by Brain Pool Plus Program (No. 2021H1D3A2A03103645) and Young Researcher Program (No. 2022R1C1C1009443) through the National Research Foundation of Korea (NRF) funded by the Ministry of Science and ICT.
Appendix A Blinking Color Balls Benchmark
A.1 Motivation
@@ A.2 Dataset Design
A.3 Challenges and Qualitative Comparison
The Blinking Color Balls tasks pose significant challenges for the models, as they are required to associate partial views of objects from temporally distant patches to learn their movement patterns and color assignment rules. We can define two levels of challenges: (1) identifying each object from the patches and predicting their positions based on their dynamics, and (2) determining the final color assignment of each object based on the given rules. The first challenge is relatively straightforward, as it primarily involves the dynamics of objects from the immediate past two frames prior to the target time step. However, the second challenge is particularly difficult, as it requires the model to reason over the entire input sequence, necessitating the identification of an object’s history from partially observed patches in a long-range context.
@@ A.3 Challenges and Qualitative Comparison
Appendix B Additional Implementation Details
B.1 SlotSSMs and OC-SlotSSMs
@@ B.2 Baseline Models
B.3 Blinking Color Balls Experiemtns
@@ B.4 Unsupervised Object-Centric Learning Experiments
B.5 3D Visual Reasoning Experiments
Model Configuration For both SlotSSM and OC-SlotSSM, we employ architectures similar to those described in Table 2, with a few modifications to accommodate the complexity of the 3D visual reasoning task. We use 12 slots and a hidden size of 64 for all modules to reduce memory consumption. Additionally, we utilize a spatial broadcast decoder for all models during the pre-training phase. Due to limited computational resources, we apply a model size of 8 for both pre-training and fine-tuning/direct training phases. It is worth noting that the model performance could potentially be further improved with a properly tuned batch size during training.
Data Augmentation To prevent overfitting, we apply data augmentation on all training phases, following the practices of prior works [44, 36]. During the model pre-training phase, we randomly sample 32 frames from the original 300 frames. Note that the sampled frames may not be consecutive, which is an intentional aspect of the data augmentation. For the fine-tuning or direct training phase, we first split the 300 frames with non-overlapping segments of length 6, then we randomly select 1 frame from each segment to construct a input sequence of length 50.
Predictor To predict the location of the snitch, we employ a transformer-based predictor. Specifically, we concatenate the learned slot representations with a learned CLS token and feed the resulting sequence into a 2-layer transformer. The output CLS token is then passed through a linear layer to predict the logits for snitch localization.
Appendix C Emerging Modularity in SlotSSMs

### v6

@@ Acknowledgements
This research was supported by GRDC (Global Research Development Center) Cooperative Hub Program (RS-2024-00436165) and Brain Pool Plus Program (No. 2021H1D3A2A03103645) through the National Research Foundation of Korea (NRF) funded by the Ministry of Science and ICT.
Appendix A Limitations & Broader Impact
Limitations SlotSSMs’ success illustrates the importance of designing architectures that align with the problem domain’s underlying modular structure. It also paves the way for future research in modular and object-centric sequence modeling. However, it has some limitations that future studies could address. First, compared to Transformer architectures, we find that SlotSSMs could benefit more from a pre-training phase in visual reasoning tasks. For example, in the 3D visual reasoning task, SlotSSMs underperform Transformer models when trained without pretraining. However, when combined with task-free pretraining, SlotSSMs demonstrate significant improvement, enabling them to outperform Transformer models. We note that this effect of task-free pre-training is more prominent in SlotSSMs than in Transformer baselines. This suggests that for tasks with sparse training signals, the sequential nature of SlotSSM performs better with a pre-training phase to learn to effectively utilize information from all time steps. We believe this phenomenon is worth further investigation in future research. Second, while the proposed architecture is applicable beyond video modeling—for example, to other modalities such as text and audio—this study has not explored these possibilities. It remains a matter for future work. Thrid, due to our academic research lab’s computing resource constraints, we were unable to significantly scale up the proposed model to industry-scale in terms of model size and data size. Scaling up SlotSSMs could uncover additional properties or limitations that are not evident at the current scale of experimentation. Lastly, future studies should investigate the effect of increased visual complexity in videos. As an initial step, we present a preliminary exploration in Appendix D.2, where SlotSSMs are applied to natural video scenes. These experiments illustrate how modularity can emerge through the independent mechanisms of SlotSSMs in real-world scenarios. We hope these findings will inspire future research on the industry-scale adoption of SlotSSMs.
Impact Statement The introduction of SlotSSMs, a novel framework that incorporates independent mechanisms into State Space Models (SSMs), has the potential to significantly impact the field of sequence modeling. By leveraging the modular structure inherent in many real-world processes, SlotSSMs offers a more intuitive and effective approach to modeling long-range temporal dependencies in object-centric video understanding and prediction tasks. The substantial performance gains demonstrated by SlotSSMs over existing sequence modeling methods highlight the importance of designing architectures that align with the underlying structure of the problem domain. This breakthrough could lead to the development of more efficient and accurate models for a wide range of applications, such as robotics, autonomous vehicles, and video surveillance systems. Moreover, the success of SlotSSMs in capturing the modular nature of real-world processes could inspire further research into modular and object-centric sequence modeling. This could result in the development of even more advanced architectures that can better handle the complexity and diversity of real-world data. Because this is a general backbone architecture for sequence modeling, it doesn’t raise direct ethical concerns. However, its ethical implications depend on the way downstream application developers use the model.
Appendix B Blinking Color Balls Benchmark
B.1 Motivation
@@ B.2 Dataset Design
B.3 Challenges and Qualitative Comparison
The Blinking Color Balls tasks pose significant challenges for the models, as they are required to reason about the object movement and color assignment rules from partial views of objects in temporally distant image patches. We can define two levels of challenges: (1) identifying the objects from image patches and predicting their future positions based on their dynamics, and (2) determining the final color assignment of each object based on the given rules. The first challenge is relatively simple, as it primarily involves learning the dynamics of objects from the past two frames prior to the target time step. However, the second challenge is particularly difficult, as it requires the model to reason over the entire input sequence, necessitating the identification of an object’s history from partially observed patches in a long-range context.
@@ B.3 Challenges and Qualitative Comparison
Appendix C Additional Implementation Details
C.1 SlotSSMs and OC-SlotSSMs
@@ C.2 Baseline Models
C.3 Blinking Color Balls Experiemtns
@@
C.4 Unsupervised Object-Centric Learning Experiments
Appendix D Additional Results
D.1 Emerging Modularity in SlotSSMs
@@
D.2 Real-World Videos Depth Estimation
In this additional experiment, we conduct a preliminary evaluation of SlotSSMs to assess its performance in real-world videos with increased visual complexity. Our aim is to observe how SlotSSMs utilize the modular representations to interpret and process real-world video data. Following previous works in object-centric learning [13], we evaluate this through a depth estimation task.
Datasets and Tasks. We select three datasets that represent distinct real-world application scenarios to observe the behavior of SlotSSMs across diverse contexts:
  1. 1.
    TikTok Dancing Dataset [33]: Commonly used for content creation tasks [31], this dataset comprises dynamic videos of individuals dancing with variety of movements and poses.
  2. 2.
    UT Egocentric Video Dataset [42]: Often utilized for egocentric action recognition [72], this dataset consists of first-person view videos that involve interactions with objects in the environment.
  3. 3.
    Waymo Open Dataset [59]: Primarily used for autonomous driving applications [43], this dataset includes videos captured from autonomous vehicles navigating traffic scenarios with diverse environmental conditions.
The primary task across these datasets is to estimate the depth of each pixel in the video frames. However, it is important to emphasize that our objective is not to develop depth estimation models that compete with existing specialized approaches. Instead, our main focus is to use this task, manageable with our lab resources, to showcase the emerging modularity in SlotSSMs for real-world video inputs.
Models. We compare OC-SlotSSM, which uses inverted attention in the Slot Encoder, against SAVi++ [13], an RNN-based object-centric learning method. Similar to the setting in Section 5, both models use a CNN encoder to extract input tokens, which are processed using inverted attention for OC-SlotSSM and slot attention for SAVi++ to produce slot representations. These slots are then used to reconstruct the image using a spatial broadcast decoder, with MSE loss as the training objective.
Figure 13: Emergent Scene Decomposition from Depth Estimation Tasks. Colors represent the ID of slots used for predicting each position. SlotSSM is capable of exploiting the inherent modular structure of real-world videos for efficient inference, without explicit segmentation supervision. For more examples please visit our project website.
Results. The quantitative results in Table 5 show that OC-SlotSSM consistently outperforms the SAVi++ baseline across all datasets, demonstrating its superior video modeling capabilities. More importantly, as illustrated by the attention patterns in Figure 13, unsupervised scene decomposition emerges during training. This demonstrates that SlotSSM is able to utilize the modular representations to discover and exploit the latent structure of the input to complete the task, while the SAVi++ baseline does not demonstrate the same level of emergent modularity.
Table 5: Depth Estimation MSE ($\downarrow$) on Real-World Datasets.
Model UT Egocentric Waymo TikTok
SAVi+++++ + 0.589 0.804 1.412
OC-SlotSSM (Ours) 0.464 0.653 1.180